How do I eliminate the fractions in the following problem: 1/5x - 3/5x =2 and 2/5x + 4/5x =3
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

1/5x -3/5x=2
Multiply the equation by 5x
1/5x*5x -3/5x*5x = 5x*2
1-3= 10x
10x=2
x=10/2
x=1/5
the second prolem is solved in the similar manner

Answer by ikleyn(53748)

.
How do I eliminate the fractions in the following problem:
(a) 1/(5x) - 3/(5x) = 2 and
(b) 2/(5x) + 4/(5x) = 3
~~~~~~~~~~~~~~~~~~~~~~~~


        In the post by @mananth, the answer is incorrect.
        I came to bring a correct solution.


(a) 1/(5x) - 3/(5x) = 2
Multiply the equation by 5x
((1/(5x))*5x - ((3/5x))*5x = 10x
1 - 3 = 10x
10x = -2
x = -2/10
x = -0.2

Solved correctly.

the second problem is solved in the similar manner